OpenStudy (tootsi123):

I have no clue what to do or where to start... Simplify (x^4)^2 without parenthesis.

imqwerty (imqwerty):

well we are given -\[(x^4)^2\]and we need to simplify it :D we know that \[(a^m)^n=a ^{m*n}\]so can u apply this in our question?

OpenStudy (tootsi123):

That (a^m)^n=a^m*n makes no since to me what so ever... is that the formula or something??

imqwerty (imqwerty):

yea thats an identity

OpenStudy (tootsi123):

Wait so you're saying if i multiply the two exponents (8) then the answer would be x^8...

imqwerty (imqwerty):

yes ur right :)

OpenStudy (tootsi123):

Yay!!! Thank you!!

imqwerty (imqwerty):

welcome :)

OpenStudy (tootsi123):

i was overanalyzing it... i hate it when i do that

imqwerty (imqwerty):

:)

imqwerty (imqwerty):

u r in which grade?

OpenStudy (tootsi123):

I am horrible at math but i am in the 12th grade

imqwerty (imqwerty):

nice :)

OpenStudy (tootsi123):

Yeah so far but this is only the 3rd day of school and its hard... my math anyways its algebra 2

imqwerty (imqwerty):

:D don worry :)

OpenStudy (tootsi123):

Thanks

imqwerty (imqwerty):

no prblem

OpenStudy (amistre64):

a^4 is defined as a, multiplied to itself 4 times: aaaa b^2 is defined as b, mulitplied to itself 2 time: bb let b = a^4 bb = a^4 a^4 = aaaa aaaa but aaaa aaaa is the definition of a^8

OpenStudy (tootsi123):

Thanks